Prove that for a particle starting from rest and moving with uniform acceleration the ratio is  = 4:3 t:s?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
As from the given data:
Equation of motion,
v=u+at (u=0)
=>t=v/a

v2=u2+2as
=>v2=2as
=>s=v2/2a

t/s = 2/v

Unless and until we know the final velocity after time t we cant find the ratio.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)